# Expanding the Boundaries of Vision Prior Knowledge in Multi-modal Large Language Models

### v1

@@ Abstract
Does the prior knowledge of the vision encoder constrain the capability boundary of Multi-modal Large Language Models (MLLMs)? While most existing research treats MLLMs as unified systems optimized through end-to-end training, the impact of vision encoder’s prior knowledge is seldom investigated. In this work, we introduce a novel metric, $Rank_{e}$, to quantify the effect of the vision encoder’s prior knowledge on MLLM performance. Our analysis reveals a positive correlation between prior knowledge and MLLM performance. Moreover, we find that domain-specific fine-tuning using solely end-to-end visual question answering (VQA) data is insufficient—particularly for entities with low inherent visual prior knowledge. To address this issue, we propose VisPRE (Vision Prior Remediation), a two-stage training framework that explicitly incorporates prior knowledge at the vision encoder level. Experimental results demonstrate that augmenting vision encoder’s prior knowledge substantially boosts the visual understanding capabilities of MLLMs, offering a novel and effective strategy for improving performance, especially in scenarios involving uncommon visual entities.
@@ 1 Introduction
Multi-modal Large Language Models have emerged as a rapidly growing area of research. Combining the powerful capabilities of Large Language Models with the ability to process visual input, MLLMs excel in tasks such as image understanding, VQA [1], image captioning, and visual instruction following. The development of models like GPT-4o[30] , GPT-4V[29], and Claude-3.5[3] have demonstrated remarkable proficiency in advanced multi-modal understanding. Besides, open-source models like LLaVA[26, 25, 18] series, Qwen2-VL[38], and InternVL2[9, 8] are making significant strides, bridging gap in the field.
@@
A pivotal challenge in advancing MLLMs is forging a seamless and robust alignment between vision and language. One effective approach involves integrating an off-the-shelf external vision encoder with a language model using a modality conversion module [2, 21, 23, 42, 13, 4, 26, 19], which we refer to as the modular approach. Compared to monolithic multi-modal approach [34, 27, 5, 41], which built from scratch using multi-modal data, modular approach is more data-efficient and achieves comparable performance. Despite these advantages, the modular approach still faces challenges, as the vision and language components are trained separately from distinct data distributions, leading to an inherent misalignment in their knowledge. To illustrate the importance of knowledge alignment, we present a knowledge quadrant diagram in Fig. 1 , with the horizontal axis representing the knowledge held by language model and the vertical axis representing the knowledge held by the vision encoder. Only when both components possess necessary knowledge (in the “Vision known & Language known” quadrant) can the multi-modal model accurately handle complex cross-modal tasks [22, 10]. Misalignment in knowledge from either vision or language side introduces limitations to the model’s capabilities, making it essential to bridge this gap to enhance the performance of multi-modal models. Many existing studies focus on addressing knowledge misalignment from the language perspective, expanding from “Vision known & Language not known” to “Vision known & Language known”. Some studies[6, 17] enhance language model knowledge with external documents related to images, while CVLM[24] trains a “Visual Knowledge Aligner” module to enrich text-based knowledge associated with images. However, as a crucial component of MLLM [12], the vision encoder also possesses varying prior knowledge about the real world, such as entities, textures and causality [32, 7]. But the impact of this vision prior knowledge on MLLM capabilities remains unexplored, leading to a natural question: How does vision prior knowledge affect MLLM’s capability? In this paper, we attempt to answer this question by investigating the following research questions:
@@
To address these questions, we introduce $Rank_{e}$ to quantify the vision encoder’s prior knowledge. Through experiments with various model combinations, we reveal a positive correlation between MLLM performance and visual prior knowledge. Fig. 2 (left) demonstrates the relationship between current MLLM performance and vision prior. Furthermore, we find that direct finetuning with end-to-end VQA data is insufficient for improving low prior entities. Fig. 2 (right) illustrates the knowledge misalignment on low prior entities. To overcome this limitation, we propose a two-stage training framework that injects vision prior knowledge into the vision encoder, resulting in significant improvements in MLLM performance. Through in-depth analysis, we demonstrate that the limitations of direct finetuning stem from the differences between next-word prediction loss and contrastive loss. In summary, our main contributions are:
@@
  • We propose a two-stage training framework VisPRE (Vision Prior Remediation) that injects prior knowledge at the vision encoder level, significantly enhancing MLLM performance, especially for entities with vision prior knowledge.
@@ 2 Vision Prior Measurement
Vision encoders are typically trained on extremely large-scale data (from 400 million to 10 billion samples[36]), often with undisclosed data (e.g. OpenAI CLIP[33]), making direct evaluation of vision priors from training data infeasible. Therefore, to answer Q1, we shift our focus to evaluating observable behavioral evidence - specifically, how effectively these encoders recognize visual entities. We thus propose the $Rank_{e}$ metric, which quantifies an encoder’s vision prior knowledge for a given entity $e$.
In this section, we begin by describing the modality alignment process in modular MLLMs, then formulating a definition of vision prior knowledge. Finally, we introduce the $Rank_{e}$ metric to quantify this knowledge.
Modern modular MLLMs establish cross-modal understanding through an alignment process that maps visual information to textual representations. Formally, given an input text prompt $T_{A}$ and target image $I_{B}$, the alignment function $\mathcal{F}$ of an MLLM generates textual description that approximate the image content:
@@
Here, $\hat{T}_{B}$ represents the generated text that preserves the semantic content of $I_{B}$. Building upon the Platonic representation hypothesis [16], we posit that cross-modal alignment occurs through a shared latent space $\mathcal{Z}$. This allows us to decompose the $P(T|I_{B})$ as:
@@
To quantify the inherent vision prior $P_{\text{vision}}(z|I_{B})$, we discretize the continuous latent space $\mathcal{Z}$ into a set of entity-specific latent representations. For a given image $I_{B}$, we approximate $P(z|I_{B})$ by evaluating the probability that the vision encoder correctly identifies entity within $I_{B}$. To achieve this, we propose the $Rank_{e}$ metric, which measures how well the encoder identifies a target entity $e$ from visual inputs, thereby evaluating the vision encoder’s inherent prior knowledge. As shown in Fig. 3, for an entity $e$, $Rank_{e}$ is computed as follows:
  1. 1.
@@ 3 Experiments
In this section, we explore the three proposed research questions. In Section 3.1, we describe the overall experimental setup. In Section 3.2, we verify the relationship between MLLM and the prior knowledge of its vision encoder. From Section 3.3 to Section 3.4, we show the insufficient of end-to-end finetuning and propose a training framework to transcend vision prior knowledge limits.
@@ 3.1 Experiment Setting
Models. To systematically examine the impact of the vision encoder’s prior knowledge on MLLM performance across different vision encoders and base LLM combinations, we train nine MLLMs from scratch based on an encoder-projector-LLM architecture. For the vision encoder, we utilize widely adopted encoders in mainstream MLLMs, including OpenAI ViT-L-14[33], SigLIP ViT-SO-14[40], and DFN ViT-H-14[15]. For the base LLM, we select the LLaVA-1.5 language model, Vicuna-7B-v1.5[11], and the recent open-source models, Llama-3.1-Instruct-7B[14] and Qwen-2.5-Instruct-7B[35]. This results in a total of nine encoder-language model combinations.
Datasets. To evaluate MLLMs under different vision priors, we require a VQA dataset that meets two conditions: (1) it provides entity annotations covering a wide range of prior knowledge—from extremely rare to very common entities; (2) it includes entity-centric visual questions and answers for MLLM performance assessment. Here, rare entities refer to those that appear infrequently or not at all in the vision encoder’s training data, making them difficult for the vision encoder to recognize accurately. The Encyclopedia-VQA[28] dataset fulfills both requirements. With extensive entity annotations covering up to 16.7k entity categories, it captures both common and rare entities and poses a hard challenge for MLLMs with its knowledge-based VQA questions.
Training. We trained the model on 8×A800 GPUs. We follow the LLaVA[26] training process and made a few modifications. We first pre-train model on LLaVA pretrain data to obtain a projector aligned with selected vision encoder. To reduce computational cost, we then sampled 10% of the LLaVA instruction tuning dataset and used it in the finetuning strategy for our subsequent experiments.
Metrics and Evaluation. We use Llama-3.1-70B[14] to judge model responses, denoted as a function $g(\cdot)$ that takes the question, entity, ground truth answer, and model output as input, returning true if the answer is correct. Using this, we define entity accuracy $\text{Acc}_{e}$ for each entity $e$ as the fraction of correct responses among all related questions:
[…]
@@
Figure 4: MLLM Performance distribution across different Ranke intervals. Performance of all MLLMs decreases as $Rank_{e}$ increases across three encoder configurations. The Vicuna-CLIP model shows an 87% performance drop from $50000<Rank_{e}<$ 5000 to $Ranke>3000$, indicating correlation between performance and vision prior knowledge. This relationship is non-linear with a critical threshold. We marked this threshold by a vertical line in the figure—green on the left indicating sufficient prior knowledge for reasoning, and red on the right showing insufficient knowledge causing sharp performance decline.
@@ 3.2 Vision Prior Constrains MLLM Performance
To investigate Q2: the impact of visual priors on MLLMs, we first categorize entities into two types: those “vision encoder knows” and those “vision encoder doesn’t know” then observe MLLM performance across both categories. Through our proposed $Rank_{e}$ metric, we measure the vision encoder’s knowledge of entities in Encyclopedia-VQA, where a lower $Rank_{e}$ indicates greater knowledge. For MLLM performance, we test accuracy in answering entity-related questions in Encyclopedia-VQA.
Our study aims at addressing knowledge misalignment where MLLM capabilities are limited by the vision encoder. Therefore, we retain only cases where the LLM component possesses adequate entity knowledge, regardless of the vision encoder’s knowledge. Specifically, we prompt the MLLM with “This is {entity name}” rather than the actual image; if the MLLM answers correctly, we retain this case. Additionally, we discovered a number of cases where MLLMs provide correct answer without image description or actual image. We attribute this to the MLLM’s dependency on question format [17]. We eliminated this subset from our analysis. Fig. 4 illustrates the relationship between MLLM accuracy and $Rank_{e}$
@@
Notably, CLIP-Vicuna MLLM does not exhibit a significant performance decline until $Rank_{e}$ reaches 3000. The phenomenon is also observed in the SigLIP and DFN configurations. This threshold effect suggests that the positive correlation between vision prior knowledge and MLLM performance is not strictly linear, but rather exhibits a mutation beyond a critical point. We posit that this stems from the vision encoder holding a known status for entities below a certain $Rank_{e}$ threshold, meaning it can still provide sufficient prior knowledge for the MLLM to answer entity-related questions. Once $Rank_{e}$ exceeds this threshold, the vision encoder no longer provides adequate prior knowledge, resulting in a sharp drop in MLLM performance. Considering that LLM part of MLLM possesses adequate knowledge about all entities here, so it is the vision encoder of MLLM that constrains the overall performance on entities beyond the threshold.
3.3 Limitations of End-to-end Finetuning
To investigate Q3: How to transcend vision prior knowledge limits, we implement a typical solution as our baseline—finetuning MLLMs on end-to-end domain-specific VQA data. Following established MLLM finetuning approaches [26, 25], we freeze the vision encoder parameters and only tune the LLM component. This setup enables the LLM parameters to compensate for limitations in vision prior knowledge.
@@ 3.3 Limitations of End-to-end Finetuning
Table 1: Dataset Statistics. We report the number of (question, answer) pairs entities for dataset split for different encoder-language model combinations. Note that the training set and testset of each dataset contain the same entities.
We construct our finetuning dataset from Encyclopedia-VQA. Following method in Section 3.2, we retain questions that MLLMs answer correctly when prompted with “This is {entity name}” instead of the actual image. After calculating $Rank_{e}$ across the dataset, we observe significantly different $Rank_{e}$ distributions for different encoders. To mitigate this distributional variance, we sample entities with more uniform rank distributions for validation. Each subset is then divided into training and test sets, ensuring they contain the same entities with no question duplication. Dataset statistics are presented in Tab. 1, with detailed construction methodology in Appendix A.
Successful knowledge-based VQA requires three MLLM capabilities: (1) recognizing entities in images; (2) possessing relevant knowledge about these entities; and (3) utilizing this knowledge to answer questions. Since we have already ensured the LLM component possesses adequate entity knowledge, MLLM performance can be enhanced through two different approaches: (1) improving entity recognition and (2) optimizing knowledge utilization.
@@
Figure 5: Perception-tuning and Knowledge-tuning performance improvement compared to Zero-shot. Perception-tuning provides limited improvement to model performance across different $Rank_{e}$ values, with a significant drop observed in Qwen-2.5’s performance within the range of $Rank_{e}>3000$. For Knowledge-tuning, we observe that in the range $Rank_{e}>3000$, both Qwen-2.5 and Llama-3.1 models finetuned with Knowledge-tuning data experience a notable decline in performance.
Finding 2: Domain-specific finetuning with only end-to-end VQA data is insufficient, particularly for entities with low visual prior knowledge. Fig. 5 illustrates the accuracy improvements of Perception-tuning and Knowledge-tuning models compared to Zero-shot baselines under CLIP encoder configuration. As shown in figure (a), after Perception-tuning, Qwen-2.5 performance decreased in the $Rank_{e}>3000$ range, while Vicuna and Llama-3.1 showed no improvement. As shown in figure (b), after Knowledge-tuning, Qwen-2.5 and Llama3.1’s performance decreased for apporoximately 33% in the $Rank_{e}>3000$ range compared to Zero-shot. The comprehensive experimental results across all nine encoder-language model combinations are shown in Tab. 2.
Figure 6: Overview of our proposed VisPRE framework. Our framework enriches the vision encoder with entity-specific prior knowledge by first generating (image, caption) pairs from entity text and then finetuning the vision encoder using contrastive loss. The enhanced encoder is subsequently integrated into the MLLM, which is further fine-tuned on Knowledgetuning data.

### v2

@@ Abstract
Does the prior knowledge of the vision encoder constrain the capability boundary of Multi-modal Large Language Models (MLLMs)? While most existing research treats MLLMs as unified systems optimized through end-to-end training, the impact of vision encoder’s prior knowledge is seldom investigated. In this work, we introduce a novel metric, $Rank_{e}$, to quantify the effect of prior knowledge of the vision encoder on MLLM performance. Our analysis reveals a positive correlation between prior knowledge and MLLM performance. Moreover, we find that domain-specific fine-tuning using solely end-to-end visual question answering (VQA) data is insufficient, particularly for entities with low inherent visual prior knowledge. To address this issue, we propose VisPRE (Vision Prior Remediation), a two-stage training framework that explicitly incorporates prior knowledge at the vision encoder level. Experimental results demonstrate that augmenting vision encoder’s prior knowledge substantially boosts the visual understanding capabilities of MLLMs, offering a novel and effective strategy for improving performance, especially in scenarios involving uncommon visual entities.
@@
Expanding the Boundaries of Vision Prior Knowledge
in Multi-modal Large Language Models
Qiao Liang1,2*, Yanjiang Liu1,2*, Weixiang Zhou2, Ben He1,2, Yaojie Lu2, Hongyu Lin2, Jia Zheng2, Xianpei Han2, Le Sun2, Yingfei Sun1 1University of Chinese Academy of Sciences 2Chinese Information Processing Laboratory, Institute of Software, Chinese Academy of Sciences {liangqiao2022,liuyanjiang2021,weixiang,luyaojie,hongyu,zhengjia}@iscas.ac.cn {benhe,yfsun}@ucas.ac.cn, {xianpei,sunle}@iscas.ac.cn
@@ 1 Introduction
Multi-modal Large Language Models have emerged as a rapidly growing area of research. Combining the powerful capabilities of Large Language Models with the ability to process visual input, MLLMs excel in tasks such as image understanding, VQA  Agrawal et al. (2016), image captioning, and visual instruction following. The development of models such as GPT-4o OpenAI (2024), GPT-4V OpenAI (2023), and Claude-3.5 Anthropic (2024) have demonstrated remarkable proficiency in advanced multi-modal understanding. Besides, open-source models like LLaVA Liu et al. (2024b, a); Li et al. (2024a) series, Qwen2-VL Wang et al. (2024), and InternVL2 Chen et al. (2024b, a) are making significant strides, bridging the gap in the field.
@@
A pivotal challenge in advancing MLLMs is forging a seamless and robust alignment between vision and language. One effective approach involves integrating an off-the-shelf external vision encoder with a language model using a modality conversion module  Alayrac et al. (2022); Li et al. (2023a, d); Zhu et al. (2023); Dai et al. (2023); Bai et al. (2023); Liu et al. (2024b); Li et al. (2022), which we refer to as the modular approach. Compared to the monolithic multi-modal approach Team (2024a); Luo et al. (2024); Bavishi et al. (2023); Zhan et al. (2024), which is built from scratch using multi-modal data, the modular approach is more data-efficient and achieves comparable performance. Despite these advantages, the modular approach still faces challenges, as the vision and language components are trained separately from distinct data distributions, leading to an inherent misalignment in their knowledge. To illustrate the importance of knowledge alignment, we present a knowledge quadrant diagram in Fig. 1, with the horizontal axis representing the knowledge held by the language model and the vertical axis representing the knowledge held by the vision encoder. Only when both components possess necessary knowledge (in the “Vision known & Language known” quadrant) can the multi-modal model accurately handle complex cross-modal tasks  Li et al. (2023c); Cheng et al. (2024). Misalignment in knowledge from either the vision or language side introduces limitations to the model’s capabilities, making it essential to bridge this gap to enhance the performance of multi-modal models. Many existing studies focus on addressing knowledge misalignment from the language perspective, expanding from “Vision known & Language not known” to “Vision known & Language known”. Some studies (Caffagni et al., 2024; Jiang et al., 2024) enhance language model knowledge with external documents related to images, while CVLM Li et al. (2024b) trains a “Visual Knowledge Aligner” module to enrich text-based knowledge associated with images. However, as a crucial component of MLLM  Collins and Olson (2014), the vision encoder also possesses varying prior knowledge about the real world, such as entities, textures, and causality  Pinker (1984); Cavanagh (2011). But the impact of this vision prior knowledge on MLLM capabilities remains unexplored, leading to a natural question: How does vision prior knowledge affect MLLM’s capability? In this paper, we attempt to answer this question by investigating the following research questions:
@@
To address these questions, we introduce $Rank_{e}$ to quantify the vision encoder’s prior knowledge. Through experiments with various model combinations, we reveal a positive correlation between MLLM performance and visual prior knowledge. Fig. 2 (left) demonstrates the relationship between current MLLM performance and vision prior. Furthermore, we find that direct fine-tuning with end-to-end VQA data is insufficient for improving MLLM performance on low prior entities. Fig. 2 (right) illustrates the knowledge misalignment on low prior entities. To overcome this limitation, we propose a two-stage training framework that injects vision prior knowledge into the vision encoder, resulting in significant improvements in MLLM performance. In summary, our main contributions are:
@@
  • We propose a two-stage training framework VisPRE (Vision Prior Remediation) that injects prior knowledge at the vision encoder level, significantly enhancing MLLM performance, especially for entities with low vision prior knowledge.
@@ 2 Vision Prior Measurement
Vision encoders are typically trained on extremely large-scale data (from 400 million to 10 billion samples Tong et al. (2024a)), often with undisclosed data (e.g., OpenAI CLIP Radford et al. (2021)), making direct evaluation of vision priors from training data infeasible. Therefore, to answer Q1, we shift our focus to evaluating observable behavioral evidence - specifically, how effectively these encoders recognize visual entities. We thus propose the $Rank_{e}$ metric, which quantifies an encoder’s vision prior knowledge for a given entity $e$.
In this section, we begin by describing the modality alignment process in modular MLLMs, then formulating the definition of vision prior knowledge. Finally, we introduce the $Rank_{e}$ metric to quantify this knowledge.
Modular MLLMs establish cross-modal understanding through an alignment process that maps visual information to textual representations. Formally, given an input text prompt $T_{A}$ and target image $I_{B}$, where $\mathcal{F}$ represents the MLLM’s internal representation function that maps inputs to hidden states, the alignment process can be described as:
@@
Here, $\hat{T}_{B}$ represents the generated text that preserves the semantic content of $I_{B}$. Building upon the Platonic representation hypothesis Huh et al. (2024), we posit that cross-modal alignment occurs through a shared latent space $\mathcal{Z}$. This allows us to decompose the $P(T|I_{B})$ as:
@@
To quantify the inherent vision prior $P_{\text{vision}}(z|I_{B})$, we discretize the continuous latent space $\mathcal{Z}$ into a set of entity-specific latent representations. For a given image $I_{B}$, we approximate $P(z|I_{B})$ by evaluating the probability that the vision encoder correctly identifies an entity within $I_{B}$. To achieve this, we propose the $Rank_{e}$ metric, which measures how well the encoder identifies a target entity $e$ from visual inputs, thereby evaluating the vision encoder’s inherent prior knowledge. As shown in Fig. 3, for an entity $e$, $Rank_{e}$ is computed as follows:
@@ 3 Experiments
In this section, we explore the three proposed research questions. In Section 3.1, we describe the overall experimental setup. In Section 3.2, we verify the relationship between MLLM and the prior knowledge of its vision encoder. From Section 3.3 to Section 3.4, we show the insufficiency of end-to-end fine-tuning and propose a training framework to transcend vision prior knowledge limits.
@@ 3.1 Experiment Setting
Models. To systematically examine the impact of vision encoder’s prior knowledge on MLLM performance across different vision encoders and base LLM combinations, we train nine MLLMs from scratch based on an encoder-projector-LLM architecture. For the vision encoder, we use widely adopted encoders in MLLMs, including OpenAI ViT-L-14 Radford et al. (2021), SigLIP ViT-SO-14 Zhai et al. (2023), and DFN ViT-H-14 Fang et al. (2023). For base LLM, we select the LLaVA-1.5 language model, Vicuna-7B-v1.5 Chiang et al. (2023), and recent open-source models, Llama-3.1-Instruct-7B Dubey et al. (2024) and Qwen-2.5-Instruct-7B Team (2024b).
Datasets. To evaluate MLLMs under different vision priors, we require a VQA dataset that meets two conditions: (1) it provides entity annotations covering a wide range of prior knowledge—from extremely rare to very common entities; (2) it includes entity-centric visual questions and answers for MLLM performance assessment. Here, rare entities refer to those that appear infrequently or not at all in the vision encoder’s training data, making them difficult for the vision encoder to recognize accurately. The Encyclopedia-VQA Mensink et al. (2023) dataset fulfills both requirements. With extensive entity annotations covering up to 16.7k entity categories, it captures both common and rare entities and poses a hard challenge for MLLMs with its knowledge-based VQA questions.
Training. We conducted training on a 8×A800 GPUs. Initially, we pre-trained the model on the LLaVA Liu et al. (2024b) dataset to develop an MLP projector aligned with selected vision encoder. For fine-tuning phase, we sampled 10% of the LLaVA instruction tuning dataset and integrated it with additional fine-tuning data to optimize computational efficiency while maintaining performance quality.
Metrics and Evaluation. We use Llama-3.1-70B Dubey et al. (2024) to judge model responses, denoted as a function $g(\cdot)$ that takes the question, entity, ground truth answer, and model output as input, returning true if the answer is correct. Using this, we define entity accuracy $\text{Acc}_{e}$ for each entity $e$ as the fraction of correct responses among all related questions:
$\text{Acc}_{e}=\frac{1}{N_{e}}\sum_{i=1}^{N_{e}}\mathbbm{1}\left[g(y_{i},\hat{y}_{i})=\text{true}\right]$ (4)
@@
Figure 4: MLLM Performance distribution across different Ranke intervals. Performance of all MLLMs decreases as $Rank_{e}$ increases across three encoder configurations. The Vicuna-CLIP model shows an 87% performance drop from $5000<Rank_{e}<$ 500 to $Rank_{e}>3000$, indicating correlation between performance and vision prior knowledge. This relationship is non-linear with a critical threshold. We marked this threshold by a vertical line in the figure—green on the left indicating sufficient prior knowledge for reasoning, and red on the right showing insufficient knowledge causing sharp performance decline.
@@ 3.2 Vision Prior Constrains MLLM Performance
To investigate Q2: “Does vision prior knowledge constrain MLLM?”, we first categorize entities into two types: those “vision encoder knows” and those “vision encoder doesn’t know” then observe MLLM performance across both categories. Through our proposed $Rank_{e}$ metric, we measure the vision encoder’s knowledge of entities in Encyclopedia-VQA, where a lower $Rank_{e}$ indicates greater knowledge. For MLLM performance, we test accuracy in answering entity-related questions in Encyclopedia-VQA.
Our study aims to address knowledge misalignment where MLLM capabilities are limited by the vision encoder. Therefore, we retain only cases where the LLM component possesses adequate entity knowledge, regardless of the vision encoder’s knowledge. Specifically, we prompt the MLLM with “This is {entity name}” rather than the actual image; if the MLLM answers correctly, we retain this case. Additionally, we discovered a number of cases where MLLMs provide correct answer without image description or actual image. We attribute this to the MLLM’s dependency on question format Jiang et al. (2024). We eliminated this subset from our analysis. Fig. 4 illustrates the relationship between MLLM accuracy and $Rank_{e}$
@@
Notably, CLIP-Vicuna MLLM does not exhibit a significant performance decline until $Rank_{e}$ reaches 3000. The phenomenon is also observed in the SigLIP and DFN configurations. This threshold effect suggests that the positive correlation between vision prior knowledge and MLLM performance is not strictly linear, but rather exhibits a mutation beyond a critical point. We posit that this stems from the vision encoder holding a known status for entities below a certain $Rank_{e}$ threshold, meaning it can still provide sufficient prior knowledge for the MLLM to answer entity-related questions. Once $Rank_{e}$ exceeds this threshold, the vision encoder no longer provides adequate prior knowledge, resulting in a sharp drop in MLLM performance. Considering that LLM part of MLLM possesses adequate knowledge about all entities here, it is the vision encoder of MLLM that constrains the overall performance on entities beyond the threshold.
3.3 Shortcomings of End-to-end Finetuning
To investigate Q3: “How to transcend vision prior knowledge limits?”, we implement a typical solution as our baseline—finetuning MLLMs on end-to-end domain-specific VQA data. Following established MLLM finetuning approaches Liu et al. (2024b, a), we freeze the vision encoder parameters and only tune the LLM component. This setup enables the LLM parameters to compensate for limitations in vision prior knowledge.
@@ 3.3 Shortcomings of End-to-end Finetuning
Table 1: Dataset Statistics. We report the number of (question, answer) pairs for each dataset split across different encoder-language model combinations. Each corresponding train-test pair shares the same entities.
We constructed our finetuning dataset from Encyclopedia-VQA. Following the method in Section 3.2, we retained questions that MLLMs answered correctly when prompted with “This is {entity_name}” instead of the actual image. After calculating $Rank_{e}$ across the dataset, we observed naturally different $Rank_{e}$ distributions across encoders. To balance the distribution of entities with varying levels of prior knowledge, we sampled entities to create more uniform rank distributions for validation. We then divided each subset into training and test sets containing the same entities but with different questions. Dataset statistics are presented in Table 1, with detailed construction methodology in Appendix A.
Successful knowledge-based VQA requires three essential MLLM capabilities: (1) recognizing entities in images; (2) possessing relevant knowledge about these entities; and (3) utilizing this knowledge to answer questions. As the LLM component already contains adequate entity knowledge, MLLM performance can be enhanced through two approaches: (1) improving visual entity recognition and (2) optimizing knowledge utilization for question answering.
@@
Figure 5: Perception-tuning and Knowledge-tuning underperform on low-prior (high Ranke) entities. The figure illustrates performance improvements compared to Zero-shot: Perception-tuning shows a significant drop for Qwen-2.5 when $Rank_{e}>3000$. Similarly, Knowledge-tuning leads to notable performance declines for both Qwen-2.5 and Llama-3.1 in the low-prior range ($Rank_{e}>3000$).
Finding 2: Domain-specific finetuning with only end-to-end VQA data is insufficient, particularly for entities with low visual prior knowledge. Fig. 5 illustrates the accuracy improvements of Perception-tuning and Knowledge-tuning models compared to Zero-shot baselines under CLIP encoder configuration. As shown in Figure (a), after Perception-tuning, Qwen-2.5 performance decreased in the $Rank_{e}>3000$ range, while Vicuna and Llama-3.1 showed no improvement. As shown in Figure (b), after Knowledge-tuning, Qwen-2.5 and Llama3.1’s performance decreased for approximately 33% in the $Rank_{e}>3000$ range compared to Zero-shot. The comprehensive experimental results across all nine encoder-language model combinations are shown in Table 2.
Figure 6: Overview of our proposed VisPRE framework. Our framework enriches the vision encoder with entity-specific prior knowledge by first extracting (image, entity_name) pairs from Perception-tuning data and then finetuning the vision encoder using contrastive loss. The enhanced encoder is subsequently integrated into the MLLM, which is further fine-tuned on Knowledge-tuning data.